The hypotenuse of a right triangle is 9 m long. The longer leg is 1 m longer than the shorter leg. Find the side lengths of the triangle.
Answer by Alan3354(69443)

The hypotenuse of a right triangle is 9 m long. The longer leg is 1 m longer than the shorter leg. Find the side lengths of the triangle.
-------------
a^2 + b^2 = 9^2
a^2 + (a+1)^2 = 81
2a^2 + 2a + 1 = 81
a^2 + a = 40
a^2 + a + 0.25 = 40.25
(a+0.5)^2 = sqrt(40.25)
a + 1/2 = sqrt(40.25)
---
a + 1/2 = -sqrt(40.25) ---> negative value, ignore